Using reasonable estimates of height differences, calculate the blood pressure difference between each of the following and the heart (specifically, pressure at the body part minus pressure at the heart): head, hand, foot. Write the answers in Pa and in mm Hg. Assuming that the pressure at the heart is 120 mm Hg, calculate the pressure at the head, at the hand, and at the foot.
Answer #1

Given,

The pressure at the heart = 120 mmHg

Let the height of the person be 180 cm

The heart is situated from head at a distance of = 60 cm

The foot is situated from heart at a distance of = 120 cm

Let the air density be 1.225 kg/m3

The pressure at head will be lesser by = 0.6 * 1.225 / 13.6

= 0.054m of Hg

= 54mm

The pressure at foot will be more by = 1.2*1.225 / 13.6

  = 0.108m of Hg

= 108 mm

P heart - P head = 54 mm of Hg

=105 * 54 / 760

=7105 Pa

P heart - P foot = 108 mm of Hg

= 105 * 108 / 760

= -14210 Pa